Ask your own question, for FREE!
Mathematics 9 Online
OpenStudy (anonymous):

Winona rolls a ten-sided die, numbered 1 through ten. What is the probability that she'll get an odd number or a 4? two-tenths four-tenths five-tenths six-tenths

OpenStudy (hoblos):

odd numbers between 1 and 10 are {1,3,5,7,9} so we have 5 odd numbers the probability that she'll get an odd = P(O) = 5/10 the probability that she'll get 4 = P(4) = 1/10 the probability that she'll get an odd number or a 4 = P(O U 4) = P(O)+P(4) = 5/10+1/10 = 6/10

Can't find your answer? Make a FREE account and ask your own questions, OR help others and earn volunteer hours!

Join our real-time social learning platform and learn together with your friends!
Latest Questions
Arriyanalol: i need help
1 day ago 7 Replies 1 Medal
Arriyanalol: CONGRATS to me
2 days ago 8 Replies 2 Medals
Legomaster346: What colleges are good that are near Arab Alabama?
1 week ago 5 Replies 0 Medals
Arriyanalol: i need help
1 week ago 6 Replies 2 Medals
Can't find your answer? Make a FREE account and ask your own questions, OR help others and earn volunteer hours!

Join our real-time social learning platform and learn together with your friends!